OpenStudy (happykiddo):

Are the two statements logically equivalent? Why or why not? (~p ∧ q) and (~p ∨ q)

OpenStudy (ash2326):

~p ^ q= \(\text{not p and q}\) ~p ∨q= \(\text{not p or q}\)

OpenStudy (ash2326):

@happykiddo do you get it? ^ = and ∨= or

OpenStudy (happykiddo):

yeah i got that

OpenStudy (ash2326):

Great:D Let me know if you have any doubt anywhere:D

OpenStudy (mathmagician):

if p=t and q=t, ~p^q=f^t=f ~pvq=fvt=t, so the statements are not equivalent

OpenStudy (happykiddo):

so they wouldnt be logically equivalent because there seperated by an "and" in the middle??

OpenStudy (mathmagician):

no, they are not logically equivalent, because there exist such arguments, that statements with these arguments gives different values.

OpenStudy (happykiddo):

hhmmm staarting to get ill go and look over the lesson again :) thanks

OpenStudy (anonymous):

you can check directly by writing the truth table for both and seeing if they are identical

OpenStudy (anonymous):

on the other hand, as mathmagician wrote, it would be a miracle if they were equivalent, because that would mean "and" meant the same thing as "or" which is certainly does not